Solve the following linear programming model using the simplex method:


maximize Z = 100x1 + 20x2 + 60x3

subject to

x3 smaller than or equal to 40

2x1 + 2x2 + 2x3 smaller than or equal to 100

3x1 + 5x2 smaller than or equal to 60


x1, x2, x3 bigger than or equal to 0


1
Expert's answer
2021-03-30T16:31:46-0400

Solution:

"\operatorname{Max} Z=100 x_{1}+20 x_{2}+60 x_{3}"

subject to

"\begin{aligned}
x_{3} & \leq 40 \\
2 x_{1}+2 x_{2}+2 x_{3} & \leq 100
\end{aligned}"

"3 x_{1}+5 x_{2} \quad \geq 60\\
and\ x_{1}, x_{2}, x_{3} \geq 0"

The problem is converted to canonical form by adding slack, surplus, and artificial variables as appropriate

1. As the constraint-1 is of type '≤' we should add slack variable "S_1"

2. As the constraint-2 is of type '≤' we should add slack variable "S_2"

3. As the constraint-3 is of type '≥' we should subtract surplus variable "S_3"  and add artificial variable "A_1"

After introducing slack, surplus, artificial variables

"\operatorname{Max} Z=100 x_{1}+20 x_{2}+60 x_{3}+0 S_{1}+0 S_{2}+0 S_{3}-M A_{1}"

subject to

"x_{3}+S_{1}
=40\\
2 x_{1}+2 x_{2}+2 x_{3}
S
=100\\
3 x_{1}+5 x_{2} \quad-S_{3}+A_{1}=60\\
and\ x_{1}, x_{2}, x_{3}, S_{1}, S_{2}, S_{3}, A_{1} \geq 0"


Negative minimum "Z_{j}-C_{j}" is "-5 M-20" and its column index is 2 . So, the entering variable is "x_{2}"

The minimum ratio is 12 and its row index is 3. So, the leaving basis variable is "A_{1}" .

Therefore, the pivot element is 5 .

Entering "=x_{2}" , Departing "=A_{1}" , Key Element =5

"R_{3}(\mathrm{new})=R_{3}(\mathrm{old}) \div 5"

"R_{1}( new )=R_{1}( old )"


"R_2(new)=R_2(old)-2R_3(new)"



Similarly, we have iteration 2, 3, 4, and 5 as follows:



Since all "Z_j-C_j\u22650"

Hence, optimal solution is arrived with value of variables as :

"x_1=50,x_2=0,x_3=0"

Max Z=5000